# Application Programming Interface CTU

### v1

@@
Geometric Model Predictive Path Integral for Agile UAV Control
with Online Collision Avoidance
Pavel Pochobradský, Ondřej Procházka, Robert Pěnička, Vojtěch Vonásek, and Martin Saska The authors are with the Multi-robot Systems Group, Faculty of Electrical Engineering, Czech Technical University in Prague, Czech Republic (http://mrs.felk.cvut.cz/). This work has been supported by the Czech Science Foundation (GAČR) under research project No. 23-06162M, by the European Union under the project Robotics and advanced industrial production (reg. no. CZ.02.01.01/00/22_008/0004590), and by CTU grant no SGS23/177/OHK3/3T/13.
@@ Abstract
In this letter, we introduce Geometric Model Predictive Path Integral (GMPPI), a sampling-based controller capable of tracking agile trajectories while avoiding obstacles. In each iteration, GMPPI generates a large number of candidate rollout trajectories and then averages them to create a nominal control to be followed by the Unmanned Aerial Vehicle (UAV). We propose using geometric SE(3) control to generate part of the rollout trajectories, significantly increasing precision in agile flight. Furthermore, we introduce varying rollout simulation time step length and dynamic cost and noise parameters, vastly improving tracking performance of smooth and low-speed trajectories over an existing Model Predictive Path Integral (MPPI) implementation. Finally, we propose an integration of GMPPI with a stereo depth camera, enabling online obstacle avoidance at high speeds, a crucial step towards autonomous UAV flights in complex environments. The proposed controller can track simulated agile reference trajectories with position error similar to the geometric SE(3) controller. However, the same configuration of the proposed controller can avoid obstacles in a simulated forest environment at speeds of up to $13\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, surpassing the performance of a state-of-the-art obstacle-aware planner. In real-world experiments, GMPPI retains the capability to track agile trajectories and avoids obstacles at speeds of up to $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
@@ I Introduction
Autonomous UAVs are increasingly being deployed in missions requiring navigation in unknown cluttered environments. In various scenarios, such as search and rescue [1], power line inspection [2], bridge inspection [3], aerial deliveries [4] and even digitization of historical monuments [5], UAVs might need to perform complex maneuvers in unknown environments while avoiding obstacles. Meanwhile, speed is an important criterion across use cases, allowing for better efficiency in the case of infrastructure inspection and even potentially helping save lives in the case of search and rescue operations. Control and local trajectory planning for UAVs in such conditions is a challenging open problem. Flying at higher speeds demands collision avoidance that supports rapid, agile maneuvers, but (small) UAVs are typically constrained by depth sensor range and computational capability.
The classical solution to real-time navigation in cluttered environments is a modular approach, where an environment map is constructed from sensor input, a global planner finds an obstacle-free trajectory, followed by a controller executing the trajectory. In this modular approach, latency can accumulate throughout the system, preventing flight at higher speeds [6]. Modularity usually also prevents the planner from considering full dynamic constraints of the UAV and the controller from accounting for obstacles, further reducing the agile flight capability. Moreover, a controller that is not aware of obstacles might cause a crash even if the original planned trajectory was collision-free due to, e.g., a tracking error. Recent learning-based methods aim to bypass the modular design by learning end-to-end mapping from sensor data to control commands [6][7]. Although such methods avoid the segmentation challenges mentioned above, they require large amounts of training data and cannot be easily reused across different UAV models [8].
@@
This letter proposes a novel UAV controller which we call GMPPI. Based on the MPPI method [9], GMPPI is a sampling-based controller utilizing a Graphics Processing Unit (GPU) for parallel generation of a large number of candidate rollout trajectories in each iteration, resulting in obstacle avoidance and agile UAV control capability. To the best of our knowledge, we present the first method that combines the use of a stereo depth camera with an MPPI-based controller. To assess if a collision occurs, we project each of the GMPPI rollout trajectories onto the latest available depth image. Similar to learning-based methods [6][7], the proposed approach unifies planning and control to improve obstacle avoidance performance, but unlike learning-based methods, it does not require any training and can be reused across different UAVs by simply modifying the relevant parameters, such as UAV mass. We propose to leverage geometric SE(3) control method [10] to generate trajectory candidates. A custom UAV-tailored cost function enables both smooth and aggressive flight depending on the shape of the reference trajectory. Finally, the rollouts employ variable-length time steps, allowing the full range of the depth sensor to be utilized without introducing additional computational complexity.
In trajectory tracking performance, the proposed GMPPI controller achieved position error similar to the SE(3) geometric controller [10] and a significantly smaller error then the existing MPPI controller implementation [9]. At the same time, the GMPPI controller is also capable of collision avoidance, unlike the other controllers. Through an ablation study, we highlight the contribution of each of the GMPPI features to overall performance, namely integration of SE(3) rollouts, introduction of a dynamic time step length, and variation of the noise and cost parameters with the time step. In obstacle avoidance capability, the GMPPI exceeded the performance of the learning-based controllers [6, 7] and of the state-of-the-art obstacle-aware Bubble planner [11]. In real-world deployment, shown in Fig. 1, the proposed GMPPI is capable of avoiding obstacles at speeds of up to $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
@@ II-1 Trajectory Tracking
Quadrotor UAV trajectory tracking can be tackled by various methods, ranging from linear Proportional Integral Derivative (PID) control to Nonlinear Model Predictive Control (NMPC[12] and SE(3) control [10]. Among these, SE(3) control stands out due to its almost global exponential stability [10], enabling aggressive flight at the limit of UAV dynamics as long as the reference trajectory is feasible. Another advantage is that it is computationally lightweight, enabling fast on-board processing. The ability to track aggressive trajectories makes it a good benchmark to compare to, while its low computational overhead allows integration of parts of SE(3) [10] into the proposed controller.
@@ II-2 Obstacle Avoidance
The task of UAVs avoiding obstacles is generally handled as part of local trajectory optimization either by a local planner [11] or a controller [13], [14]. One of the oldest methods of real-time replanning and optimization of collision-free trajectories utilized Artificial Potential Field (APF[15], where an artificial force defined as the gradient of an APF acts on the trajectory, the goal position has the lowest potential, and states where a collision would occur have a high potential. Alternatively, Vector Field Histogram (VFH) [16] and VFH+ [17] discretize the possible trajectories into a polar histogram and balance path smoothness, distance to goal, and obstacle avoidance. Both APF and VFH+ have been applied to UAV obstacle avoidance in [18] and [19], respectively.
@@
However, APF and VFH+ cannot take full advantage of the agility of some UAV systems. Directly finding a safe, e.g., time-optimal trajectory utilizing the full UAV agility while avoiding obstacles is not computationally feasible in real time. Instead of directly optimizing the trajectory, precomputed motion primitives can be used to build a feasible trajectory. Such primitives can be state-based, control-based, or motion-based [20]. Improvements to motion primitive-based search algorithms for local obstacle-aware trajectory planning were developed in [21]. To further decrease trajectory computation time and therefore latency, in addition to using precomputed motion primitives, the mapping step can be bypassed. Methods such as [20] or [22] use most recent image data to generate local trajectories. Despite integration of these methods into complete navigational pipelines and extensive testing [23], they cannot utilize the full dynamic capabilities of agile UAVs.
@@
Due to the fact that the quadrotor UAV is differentially flat [24], an alternative to motion primitive-based planning is to directly solve the trajectory optimization problem by creating polynomial trajectories [25]. The method introduced in [25] has been applied along with a corridor planner in [11], with performance exceeding some methods based on Imitation Learning (IL[6]. Obstacle avoidance can also be bundled with control by using an adapted version of Model Predictive Control (MPC), where obstacle avoidance can be included as a condition of trajectory optimization [13], but it is computationally demanding.
@@
Although recent methods utilizing separate mapping, planning and control modules have significantly advanced agile flight capabilities [11], there has also been a growing interest in alternatives that mitigate latency, such as learning-based approaches that generate control commands directly from sensor input [6], using either Reinforcement Learning (RL[26] or IL [6]. Although these methods produce impressive results, they embed very little prior information about the systems they are controlling, treating them instead as black boxes. The controller must therefore be retrained after any change to the parameters of the controlled UAV, such as mass. Additionally, RL and IL methods have very low sample efficiency [27]. Recent research has suggested mitigating the low sample efficiency by using differentiable simulations [7], but this approach does not mitigate the other drawbacks of learning-based methods.
@@ II-3 Model Predictive Path Integral
MPPI is a variant of MPC utilizing principles of Path Integral [28] control. Where traditional MPC solves the local trajectory optimization problem using iterative optimization methods, MPPI uses a Monte Carlo sampling-based approach [29]. This allows MPPI to work with gradient-free and even non-smooth cost functions.
@@
More recently, Perception-Aware MPPI (PA-MPPI[32] was introduced, which augments MPPI with perception objectives for exploration when the goal is occluded. This extends MPPI toward global navigation, while our proposed approach focuses on local planning and control, enabling agile flight with collision avoidance.
@@ III Methodology
This section begins with a brief overview of the MPPI framework. Then it details the architecture of the proposed GMPPI controller, highlighting the integration of depth sensing, implementation of rollouts generated using an SE(3) controller, and the development of a custom cost function enabling agile and smooth flight as well as obstacle avoidance.
MPPI, when used as a controller, is capable of tracking trajectories while avoiding obstacles [9]. It is first initialized with a nominal control sequence $\bm{\mathbf{u}}^{\mathrm{nom}}=\begin{bmatrix}\bm{\mathbf{u}}^{\mathrm{nom}}_{0},\ldots,\bm{\mathbf{u}}^{\mathrm{nom}}_{N-1}\end{bmatrix}$, where the lower index $j$, $0\leq j<N$, indicates a value at the $j$th time step of a sequence (i.e., $u_{j}\in\bm{\mathbf{u}}^{\mathrm{nom}}$). In each iteration, $K$ disturbance sequences $\delta{\bm{\mathbf{u}}_{j}^{k}}$, each of length $N$, are sampled from a normal distribution with zero mean. Rollout commands $\bm{\mathbf{u}}_{j}^{k}$ and states $\bm{\mathbf{x}}_{j+1}^{k}$ are then computed as
@@
$\dot{\bm{\mathbf{p}}}$ $=\bm{\mathbf{v}},$ (2)
@@
$\dot{\bm{\mathbf{q}}}$ $=\frac{1}{2}\bm{\mathbf{q}}\odot\begin{bmatrix}0\\ \bm{\mathbf{\omega}}\end{bmatrix}$,
$\dot{\bm{\mathbf{\omega}}}$ $=\mathbf{J}^{-1}\left(\bm{\mathbf{\tau}}-\bm{\mathbf{\omega}}_{c}\times\mathbf{J}\bm{\mathbf{\omega}}_{c}\right),$
@@
Each rollout state sequence $\bm{\mathbf{x}}^{k}$ is evaluated by a cost function. The proposed GMPPI cost function is presented in (17). Rollout costs $C^{k}$ are transformed into weights $w_{k}$ using
@@ III-A Depth Camera Integration
To the best of our knowledge, no implementation of MPPI has yet been integrated with a depth sensor for agile collision avoidance. Enabling obstacle avoidance during fast flight requires minimizing the delay between sensor input and control output. Instead of constructing and performing obstacle avoidance based on an environment map or processing the input image with a machine learning model [6], the availability of rollout states $\bm{\mathbf{x}}^{k}_{j}=\begin{bmatrix}{\bm{\mathbf{p}}^{k}_{j}}^{T}&{\bm{\mathbf{v}}^{k}_{j}}^{T}&{\bm{\mathbf{q}}^{k}_{j}}^{T}&{\bm{\mathbf{\omega}}^{k}_{j}}^{T}\end{bmatrix}^{T}$ is used.
@@ III-B Dynamic Rollout Time Steps
Small agile quadrotors are limited in terms of the hardware they can carry on board, which in turn limits the available computational power. This restricts both the number of GMPPI rollouts as well as the number of steps in each rollout.
To ensure that the sensor range is fully utilized at any flight speed, a vector $\bm{\mathbf{n}}=\begin{bmatrix}n_{0},\ldots,n_{N-1}\end{bmatrix}$ is defined, where each element $n_{j}$ acts as a time step multiplier at the $j$th step of each rollout. The corresponding time step is calculated as $\Delta{t}_{j}=n_{j}\Delta{t}_{0}$, where $\Delta{t}_{0}=0.01$\mathrm{s}$\$ is the base controller update period.
@@
which ensures that the range of the used sensor is fully utilized. The vector n has values

### v2

@@
Geometric Model Predictive Path Integral for Agile UAV Control with Online Collision Avoidance
Pavel Pochobradský, Ondřej Procházka, Robert Pěnička, Vojtěch Vonásek, Martin Saska Manuscript received: October, 9, 2025; Revised January, 13, 2026; Accepted February, 8, 2026.This paper was recommended for publication by Editor G. Loianno upon evaluation of the Associate Editor and Reviewers’ comments. The authors are with the Multi-robot Systems Group, Faculty of Electrical Engineering, Czech Technical University in Prague, Czech Republic (http://mrs.felk.cvut.cz/). This work has been supported by the Czech Science Foundation (GAČR) under research project No. 23-06162M, by the European Union under the project Robotics and advanced industrial production (reg. no. CZ.02.01.01/00/22_008/0004590), and by CTU grant no SGS23/177/OHK3/3T/13. Digital Object Identifier (DOI): see top of this page.
@@ Abstract
In this letter, we introduce Geometric Model Predictive Path Integral (GMPPI), a sampling-based controller capable of tracking agile trajectories while avoiding obstacles. In each iteration, GMPPI generates a large number of candidate rollout trajectories and then averages them to create a nominal control to be followed by the controlled Unmanned Aerial Vehicle (UAV). Classical Model Predictive Path Integral (MPPI) faces a trade-off between tracking precision and obstacle avoidance; high-noise random rollouts are inefficient for tracking but necessary for collision avoidance. To this end, we propose leveraging geometric SE(3) control to generate a portion of GMPPI rollouts. To maximize their benefit, we introduce a UAV-tailored cost function balancing tracking performance with obstacle avoidance. All generated rollouts are projected onto depth images for collision avoidance, representing, to our knowledge, the first method utilizing depth data directly in a UAV MPPI loop. Simulations show GMPPI matches the tracking error of an obstacle-blind geometric controller while exceeding the avoidance capabilities of state-of-the-art planners and learning-based controllers. Real-world experiments demonstrate flight at speeds up to 17 m/s and obstacle avoidance up to 10 m/s.
@@ I Introduction
Autonomous UAVs are increasingly being deployed in missions requiring navigation in unknown cluttered environments. In various scenarios, such as search and rescue [1], power line inspection [2], and even digitization of historical monuments [3], UAVs might need to perform complex maneuvers in unknown environments while avoiding obstacles. Meanwhile, speed is an important criterion across use cases, allowing for better efficiency in the case of inspection and even helping save lives in the case of search and rescue. Control and local trajectory planning for UAVs in such conditions is a challenging open problem. Flying at higher speeds demands collision avoidance that supports rapid, agile maneuvers, but (small) UAVs are typically constrained by depth sensor range and computational capability.
The classical solution to real-time navigation in cluttered environments is modular: an environment map is built from sensor input, a global planner finds an obstacle-free trajectory, and a controller executes it. In this approach, latency can accumulate, preventing flight at higher speeds [4]. Modularity often prevents the planner from considering full dynamics and the controller from accounting for obstacles, reducing agility. Moreover, a controller that is not aware of obstacles might cause a crash even if the original planned trajectory was collision-free due to, e.g., a tracking error. Recent learning-based methods aim to bypass the modular design by learning end-to-end mapping from sensor data to control commands [4][5]. Although such methods avoid the segmentation challenges mentioned above, they require large training data and cannot be easily reused for different UAVs [6].
@@
This letter proposes a novel UAV controller which we call GMPPI. Based on the MPPI method [7], GMPPI is a sampling-based controller utilizing a Graphics Processing Unit (GPU) for parallel generation of a large number of candidate rollout trajectories in each iteration, resulting in obstacle avoidance and agile UAV control capability. Like learning-based methods [4, 5], our approach unifies planning and control, but requires no training and is reusable across UAVs by modifying parameters. Tuning existing MPPI methods [7] requires balancing trajectory tracking precision, flight smoothness, and obstacle avoidance capability. In particular, rollout trajectories generated by adding noise to a nominal trajectory do not represent an efficient way of following a reference trajectory.
Therefore, in addition to random rollouts, we propose using geometric SE(3) control [8] to generate a portion of the GMPPI rollout trajectories, enabling precise reference tracking. Notably, even a small number of such SE(3)-based rollouts significantly improves trajectory tracking performance compared to prior MPPI approaches [7]. To maximize this benefit, we introduce a UAV-tailored GMPPI cost function, which enables us to retain the tracking capability of SE(3) control alongside the obstacle avoidance capability. Furthermore, to the best of our knowledge, we introduce the first method utilizing depth sensor data directly in a UAV MPPI-based controller loop. We also adopt a variable time-step sampling strategy, balancing the importance of near and distant trajectory segments while fully exploiting sensor range. Simultaneously, deterministic yaw control improves sampling performance without increasing the number of samples.
We show that in trajectory tracking tasks, the proposed GMPPI controller achieves a tracking error similar to geometric SE(3) control [8] and significantly smaller than an existing MPPI controller [7]. Through an ablation study, we highlight the contribution of each of the GMPPI features to overall performance, namely integration of SE(3) rollouts, variation of noise and cost parameters with time step and dynamic rollout time step length. In obstacle avoidance capability, GMPPI exceeded the performance of the learning-based controllers [4, 5] and of the state-of-the-art obstacle-aware Bubble planner [9]. In real-world deployment, shown in Fig. 1, the proposed controller is capable of avoiding obstacles at speeds of up to $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
@@ II-1 Trajectory Tracking
Quadrotor trajectory tracking ranges from linear Proportional Integral Derivative (PID) to Nonlinear Model Predictive Control (NMPC[10] and SE(3) control [8]. SE(3) control stands out due to its almost global exponential stability [8], enabling aggressive flight at the limit of dynamics as long as the reference trajectory is feasible. Its low computational overhead allows fast on-board processing and integration into the proposed controller, while its tracking capability makes it a suitable benchmark.
@@ II-2 Obstacle Avoidance
The task of UAVs avoiding obstacles is generally handled as part of local trajectory optimization either by a local planner [9] or a controller [11] One of the oldest methods of real-time replanning and optimization of collision-free trajectories utilized Artificial Potential Field (APF[12], where an artificial force defined as the gradient of an APF acts on the trajectory, the goal position has the lowest potential, and states with a possible collision have a high potential. Alternatively, Vector Field Histogram (VFH) [13] and VFH+ [14] discretize the possible trajectories into a polar histogram and balance path smoothness, distance to goal, and obstacle avoidance. Both APF and VFH+ have been applied to UAV obstacle avoidance in [15] and [16], respectively.
@@
APF and VFH+ cannot take full advantage of the agility of some UAVs. Directly finding a safe, e.g., time-optimal trajectory utilizing the full UAV agility while avoiding obstacles is not computationally feasible in real time. Instead of directly optimizing the trajectory, precomputed motion primitives can be used to build a feasible trajectory. Such primitives can be state-based, control-based, or motion-based [17]. Improvements to motion primitive-based search algorithms for local obstacle-aware trajectory planning were developed in [18]. To further decrease trajectory computation time and latency, the mapping step can be bypassed.
Methods such as [17] or [19] use most recent image data to generate local trajectories. Despite integration of these methods into navigational pipelines and extensive testing [20], they cannot utilize the full dynamic capabilities of agile UAVs.
@@
As the quadrotor UAV is differentially flat [21], an alternative to motion primitive-based planning is to directly solve the trajectory optimization problem by creating polynomial trajectories [22]. The method [22] has been applied along with a corridor planner in [9], with performance exceeding some methods based on Imitation Learning (IL[4]. Alternatively, obstacle avoidance can be included as a condition of trajectory optimization [11] in an adapted Model Predictive Control (MPC) controller, though this is computationally demanding.
@@
Although recent methods utilizing separate mapping, planning and control modules have significantly advanced agile flight capabilities [9], alternatives mitigating latency have been explored, such as learning-based approaches that generate control commands from sensor inputs [4], using either Reinforcement Learning (RL[23] or IL [4]. Although these methods produce impressive results, they embed very little prior information about the systems they are controlling, treating them instead as black boxes. The controller must therefore be retrained after any change to the parameters of the controlled UAV, such as mass. Additionally, RL and IL methods have very low sample efficiency [24]. Recent research has suggested mitigating the low sample efficiency by using differentiable simulations [5], but it does not mitigate the other drawbacks of learning-based methods.
@@ II-3 Model Predictive Path Integral
MPPI is a variant of MPC utilizing principles of Path Integral [25] control. While traditional MPC solves local trajectory optimization using iterative methods, MPPI uses a Monte Carlo sampling-based approach [26], allowing it to work with gradient-free and non-smooth cost functions.
@@
More recently, Perception-Aware MPPI (PA-MPPI[29] was introduced, which augments MPPI with perception objectives for exploration when the goal is occluded. This extends MPPI toward global navigation, while our approach focuses on local planning and control for agile flight with collision avoidance.
@@ III Methodology
This section begins with a brief overview of MPPI control. Then it details the architecture of the proposed GMPPI controller, highlighting the integration of depth sensing, implementation of rollouts generated using an SE(3) controller, and the development of a custom cost function enabling agile and smooth flight as well as obstacle avoidance.
The base MPPI terminology and notation shown in this section are adopted from [7], where the complete method is presented. MPPI, when used as a controller, is capable of tracking trajectories while avoiding obstacles. It is first initialized with a nominal control sequence $\bm{\mathbf{u}}^{\mathrm{nom}}=\begin{bmatrix}\bm{\mathbf{u}}^{\mathrm{nom}}_{0},\ldots,\bm{\mathbf{u}}^{\mathrm{nom}}_{N-1}\end{bmatrix}$, where the lower index $j$, $0\leq j<N$, indicates a value at the $j$th time step of a sequence (i.e., $u_{j}\in\bm{\mathbf{u}}^{\mathrm{nom}}$). In each iteration, $K$ disturbance sequences $\delta{\bm{\mathbf{u}}_{j}^{k}}$, each of length $N$, are sampled from a normal distribution with zero mean. Rollout commands $\bm{\mathbf{u}}_{j}^{k}$ and states $\bm{\mathbf{x}}_{j+1}^{k}$ are then computed as
@@
$\dot{\bm{\mathbf{p}}}$ $=\bm{\mathbf{v}},\text{ }\dot{\bm{\mathbf{q}}}=\frac{1}{2}\bm{\mathbf{q}}\odot\begin{bmatrix}0\\ \bm{\mathbf{\omega}}\end{bmatrix},\text{ }\dot{\bm{\mathbf{\omega}}}=\mathbf{J}^{-1}\left(\bm{\mathbf{\tau}}-\bm{\mathbf{\omega}}_{c}\times\mathbf{J}\bm{\mathbf{\omega}}_{c}\right),$ (2)
@@
Each rollout state sequence $\bm{\mathbf{x}}^{k}$ is evaluated by a cost function to produce a cost $C^{k}$. Trajectories colliding with an obstacle and trajectories further from the reference receive higher costs. The proposed GMPPI cost function mapping rollout state sequences $\bm{\mathbf{x}}^{k}$ to costs $C^{k}$ is presented in Sec. III-E. Costs $C^{k}$ are used to compute rollout weights $w_{k}$ using
@@ III-A Depth Camera Integration
To the best of our knowledge, no MPPI implementation has yet been integrated with a depth sensor for agile collision avoidance. To minimize delay, we avoid map construction or heavy model processing [4]. Instead, we use the available rollout states $\bm{\mathbf{x}}^{k}_{j}=\begin{bmatrix}{\bm{\mathbf{p}}^{k}_{j}}^{T}&{\bm{\mathbf{v}}^{k}_{j}}^{T}&{\bm{\mathbf{q}}^{k}_{j}}^{T}&{\bm{\mathbf{\omega}}^{k}_{j}}^{T}\end{bmatrix}^{T}$.
@@ III-B Dynamic Rollout Time Steps
Small UAVs are limited in terms of the on-board hardware, which in turn limits the computational power. This restricts the number of rollouts and the number of steps in each rollout.
To ensure that the sensor range, labelled $s$, is fully utilized at any flight speed, a vector $\bm{\mathbf{n}}=\begin{bmatrix}n_{0},\ldots,n_{N-1}\end{bmatrix}$ is defined, where each element $n_{j}$ acts as a time step multiplier at the $j$th step of each rollout. The corresponding time step is calculated as $\Delta{t}_{j}=n_{j}\Delta{t}_{0}$, where $\Delta{t}_{0}=0.01$\mathrm{s}$\$ is the base controller update period.
@@
which ensures that the range $s$ of the used sensor is fully utilized. The vector n has values
@@
Aside from the sensor usage benefits, shorter time steps at the beginning of the rollout increase the importance of near-term flight precision and thus stability of GMPPI in hover.